ed to withdraw the money for the first time. The account current y has $4,037 in it and pays a n 4% interest rate. Your grandfather put some money n an account for you on the day you were bom You are now 18 years old and are alo a. How much money would be in the account if you left the money there until your 25th birthday? b. How much would be in your account if you left the money in the account until your 65th birthday? c. How much money did your grandfather originally put in the account? a. How much money would be in the account if you left the money there until your 25th birthday? The future value is $Round to the nearest dollar.) b. How much would be in your account if you left the money in the account until your 65th birthday? The future value is (Round to the nearest dollar.) c. How much money did your grandfather originally put in the account? The present value is s(Round to the nearest dollar.)
Answer #1

a) until 25th birthday = 4037*1.04^7 = $5312.41

b) until 65th birthday = 4037*1.04^47 = $25505.02

c) inital sum = 4037/1.04^18 =$1992.77